OpenStudy (ray_masters):

can someone help me Asap plz with two questions? Which statement is true for the equation 3x − 3x − 2 = −2? It has no solution. It has one solution. It has two solutions. It has infinitely many solutions. Which statement is true about the solutions for the equation 4y + 6 = −2? It has infinitely many solutions. It has two solutions. It has one solution. It has no solution.

Mehek (mehek14):

3x − 3x − 2 = −2 Combine like terms 3x - 3x = ?

OpenStudy (ray_masters):

isn't that just 0?

Mehek (mehek14):

Yes so 0 - 2 = -2 which makes it -2 = -2 infinite solutions

OpenStudy (ray_masters):

ohh, i get that, so for the 1th one it's D

Mehek (mehek14):

Yes because no matter what value of x you put, the ending result will always be -2 = -2

Mehek (mehek14):

For the second one, there is no x?

OpenStudy (ray_masters):

and 4 the second 1, i know that 4y + 6 = -2 and that y is -2 right? and yea there's no x

Mehek (mehek14):

Right

OpenStudy (ray_masters):

does that mean there's only 1 solution?

Mehek (mehek14):

Yep

OpenStudy (ray_masters):

ok thanks for your help i get them now :)

OpenStudy (ray_masters):

do you think you could check over another question that i did?

OpenStudy (ray_masters):

Which statement is true for the equation 4n − 3 = 4n − 2? It has no solution. ~~~~ It has one solution. It has two solutions. It has infinitely many solutions.

Mehek (mehek14):

Correct

OpenStudy (ray_masters):

yay! lol thanks again for your help i got a 100% :D

Mehek (mehek14):

You're welcome :)